### v1

@@ Abstract
We investigate whether the hidden states of large language models (LLMs) can be used to estimate and impute economic and financial statistics. Focusing on county-level (e.g. unemployment) and firm-level (e.g. total assets) variables, we show that a simple linear model trained on the hidden states of open-source LLMs outperforms the models’ text outputs. This suggests that hidden states capture richer economic information than the responses of the LLMs reveal directly. A learning curve analysis indicates that only a few dozen labelled examples are sufficient for training. We also propose a transfer learning method that improves estimation accuracy without requiring any labelled data for the target variable. Finally, we demonstrate the practical utility of hidden-state representations in super-resolution and data imputation tasks.
@@ 1 Introduction
During training, generative large language models (LLMs) are exposed to vast amounts of information, including data relevant to economic modelling, such as geospatial statistics and firm-level financial metrics. If LLMs can effectively retrieve and utilise this knowledge, they could reduce dependence on external data sources that are time-consuming to access, clean, and merge, or that incur financial costs. Moreover, if LLMs accurately represent data, they could support downstream tasks like data imputation and outlier detection. In this study, we evaluate if and how LLMs can be used for typical economic data processes.
@@
We test the usefulness of LLMs for data tasks with a large array of empirical analyses on regional economic data of the US, UK, EU, and Germany and financial data on US listed firms. First, we show that a regularised linear model learned on the hidden states often provides substantially better estimates of the economic and financial variables than the text output of the model, particularly for less common statistics. This result holds across open-source LLMs of varying size (1–70 billion parameters). A learning curve analysis finds that usually small samples of labelled data suffice to learn an accurate linear model on the embeddings. We also suggest a simple transfer learning algorithm that allows the model to estimate a statistical variable without any labelled data for that variable. We achieve this by both learning from other labelled variables and using the text output of the LLMs as noisy labels on the variable of interest.
@@
Finally, we investigate the practical relevance of our findings for two data processes. Firstly, we show that we can exploit the embeddings to consistently improve the imputation of economic and financial variables, a common challenge in both research and industrial data pipelines. Secondly, we provide evidence that hidden states of LLMs can help in super-resolution tasks, where we want to estimate economic statistics for a finer-grained geographical level given data at a coarser level.
Our findings strongly support the value of LLMs in data processing tasks, which warrants attention, given that tasks such as identifying data sources, extracting numeric information, merging data series, and handling missing values are time-consuming and error-prone. This is evident from the numerous data providers that sell packaged, cleaned, and well-structured datasets, even when a large proportion of the underlying data is publicly accessible online.
This study is structured as follows. Section 1.1 discusses the relevant literature, Section 2.1 describes the dataset and Section 2 outlines our empirical approach. The main results on the accuracy of our approach to estimate economic variables from embeddings are shown in Section 3. Section 4 presents the methodology and results for two data applications: imputation and super-resolution. Section 5 concludes.
@@ 1.1 Literature review
Our study is most closely related to papers that investigate how well LLMs can recall geospatial statistics. Manvi et al. (2023) show that both GPT3.5 and fine-tuned Llama 2 can be prompted to reproduce geospatial information, such as population density, mean income, education, or house prices. The LLMs recall these variables more accurately than supervised machine learning methods, which inferred the information from other data sources such as satellite images of luminosity at night. Note that the authors only consider the text output of the LLM and do not exploit the hidden states as we do.
@@
Similarly, Li et al. (2024) test how well LLMs can reproduce a large set of variables about cities and regions such as population density, life expectancy, average travel time, and number of patents. As in our study, this paper tests the accuracy of both the text output and a linear model on the hidden layer activations to estimate the statistics. However, they focus on a third approach: they ask the LLM to identify relevant features that can help to predict a variable of interest and then ask the LLM to generate values for these variables. After feeding these variables into a supervised ML models to predict the variable of interest, this approach performs on par with a linear model on the hidden states in 5-fold cross-validation. In contrast to our work, this paper purely focuses on how well the LLMs can recall statistics and does not consider transfer learning or downstream tasks.
@@
Our main empirical approach of learning a linear model on hidden states of a language model is known as linear probing in the literature (Belinkov, 2022; Alain and Bengio, 2016) and has been used to understand the inner workings and capabilities of LLMs. Our study builds in particular on the findings by Gurnee and Tegmark (2023), showing that hidden states of LLMs represent space and time. Concretely, a linear model trained on the hidden states of Llama 2 accurately infers the death of historical figures, the release year of artworks, or the geographical location of landmarks. Godey et al. (2024) investigate the scaling laws of these results, showing that larger models represent space and time more accurately. Further, Chen et al. (2023) perturb activations in the LLM to show that there is a causal connection between the representations and the text output.
@@
Zhu et al. (2024) use linear probing to show that LLM embeddings represent the results of simple addition problems. While the exact number cannot be reconstructed using linear probes, the correlation between the actual number output and the inferred number (via the probes) is high. Linear probing is also used to gauge the truthfulness of statements (Marks and Tegmark, 2023; Orgad et al., 2024; Bürger et al., 2024) and Liu et al. (2024) show that hidden states can also be used to estimate uncertainty of LLM outputs in different tasks.
@@
Our work also relates to studies that use hidden states of LLMs for downstream tasks such as classification. Particularly, Buckmann and Hill (2024) showed how a linear model trained on hidden states of small LLMs after task-specific prompting can outperform GPT-4 when trained on only a few dozen samples per class (see also (Cho et al., 2023)). Hidden states of LLMs have also proven to perform well as general text embedding Lee et al. (2024); Rui Meng (2024) as evidenced by their competitive performance on the Massive Text Embeding Benchmark (MTEB) (Muennighoff et al., 2022), which includes tasks such as clustering, re-ranking, retrieval and classification.
Our approach to transfer learning is inspired by work on deep learning from noisy labels (see Song et al., 2022, for a review). Oyen et al. (2022) show that neural networks are surprisingly robust to label noise and Liu et al. (2020) find that deep learning models learn to predict correct labels in the early stage of learning and overfit to noisy labels (memorisation) only in later stages.
@@
The literature on the use of LLMs for data imputation and super-resolution tasks – the two practical applications of our methodology that we explore in this study – is scarce. Hayat and Hasan (2024) fine-tuned Llama 2 to impute missing values given a textual description of the instance’s features. Ding et al. (2024) use a similar approach for data imputation in recommender systems. By contrast, we do not rely on text output but use the embeddings of the LLM as additional features when employing standard imputation methods.
@@
While there exists a rich statistical and machine learning literature on the estimation of geospatial statistics at a granular level (Tzavidis et al., 2018; Chi et al., 2022; Aiken et al., 2022; Viljanen et al., 2022), we are not aware of any study that explorse the use of LLMs for this purpose.
@@ 2.1 Datasets
We use four public regional datasets covering US counties, EU regions, UK districts and German districts. Additionally, we use a dataset containing public financial information about US listed firms. The datasets are listed in Table 1. We assembled the datasets using different sources, which are described in detail in Appendix A. We provide more extensive results for our main datasets, US counties and US listed firms, which contain a larger number of observations than the other datasets.
@@
Table 1: Datasets.
Key variables that we use in the regional datasets are population, GDP per capita, unemployment rate, a measure of the income per capita, and a measure of life expectancy. Key variables we observe in the US firms dataset are total assets, market capitalisation, profitability metrics such as return on equity and return on assets. For the super-resolution analysis in Section 4.2 we use additional data that we describe in that Section.
@@ 2.2 Experimental set-up
We exclusively use open-source LLMs, as, to the best of our knowledge, none of the providers of the leading proprietary models provide access to the hidden states of their models. Furthermore, we focus on smaller models because these are easier and cheaper to use. The models we use are listed in Table 2. In the main experiments, we test three models of the Llama 3 family (1B, 8B and 70B parameters) and the Phi-3-mini (3.8B parameters) model. Apart from Llama 3 70B, we accessed the models through the transformers library (Wolf et al., 2020) on an NVIDIA Tesla V100 GPU. Due to its large size, we use a quantised version of Llama 3 70B on 2 NVIDIA T4 GPUs and obtain the embeddings with the llama-cpp-python library. To obtain the text outputs, we use the non-quantised model hosted on replicate.com.
We also consider a reasoning model: Qwen QwQ (32B parameters). To assess whether the reasoning paradigm improves the estimation of regional and financial statistics we pit the model directly against Qwen 2.5 32B, which is the general-purpose LLM on which Qwen QwQ is based. To obtain the text output of both models, we use the instances hosted on groq.com. We obtain the embeddings of Qwen 2.5 32B from a quantised version using the llama-cpp-python.
@@
When learning a linear model on the embeddings, we also test a generic prompt, which does not vary with the variable of interest but just embeds the name of the entity (e.g. “Orange County, California”).
@@ 2.3 Linear model on embeddings (LME)
We calibrate a linear model on the embeddings (LME) of the last token of the prompt. Our baseline model has 32 layers, with a 4096-dimensional embedding vector. If not stated otherwise, we use the embeddings of the 25th layer because LME performs better on this layer than on others (Figure AV in the appendix).
To estimate the performance of LME, we employ grouped cross-validation on the regional datasets. Specifically, we ensure that the training and test sets do not share any values of the grouping variable. The grouping variable for each dataset is shown in Table 1. For US counties, for example, the grouping variable is US states. Our baseline results are based on 5-fold cross-validation but we also report learning curves, where, adhering to the grouping, we sample training sets of increasing size.
We either train the linear model on the embeddings directly or on the first $k$ PCA components, with $k\in\{5,10,25,50,100,200\}$. We fit the PCA to all observations and not only on the training set.
In the following, we denote the ground truth values of the variable of interest as $y$ and the embedding matrix as $E$. The prediction model, usually a linear model, is denoted as $m$ and the predicted values $\hat{y}=m(E)$. The LLM’s text output is denoted as $\overset{txt}{y}$. In some analyses, we train a model on data of several variables or several datasets. In these cases we use subscripts $y_{v},E_{v}$ or $y_{d},E_{d}$ to respectively denote variables as $v$ and datasets as $d$.
@@ 2.4 Data transformations and performance metrics
To avoid a strong influence of outliers on the performance metrics and on the estimation of the LME, we transform some of the dependent variables. Specifically, across all datasets, we applied a log transformation to non-negative variables with a skewed distribution. For skewed variables with negative values, we use a cubic transformation ($f(x)=sgn(x)\times|x|^{\frac{1}{3}}$). The full list of variables in all datasets and the transformations applied are shown in Table LABEL:tab:tab_cv_all.
When assessing the performance of the text output and LME, we test both models on exactly the same observations. The text outputs of the LLMs sometimes fall outside the range of expected values (e.g. unemployment rate $>50$%). In this case, we decide to remove observations to reduce noise. Parsing numeric values from LLMs’ textual answers is challenging, due to the inconsistent format and inconsistent use of units. We went to considerable effort to correctly parse the text output and conducted comprehensive manual checks on all variables to ensure the accuracy of our parsing approach. Our parsing approach is described in detail in Appendix B.2.
@@ 3 Using LLM embeddings to infer statistics
Figure 1: Cross-validation performance. The variable labels are shortened. The full variable names used when querying the LLM are shown in Table LABEL:tab:tab_cv_all
@@ 3.1 Cross-validation
Figure 1 compares the performance (Spearman correlation) of the text output to the baseline LME learned on the embeddings of the completion prompt and the generic prompt. The variables are ordered by decreasing performance of the text output. LME based on the completion prompt performs better than LME on the generic prompt in all datasets except the US firms data, where both perform equally well on most variables. We therefore focus on the completion prompt in the following analyses.
LME beats the text output most of the time. The performance advantage of LME increases with the decreasing performance of the text output. On more common variables, such as population, or unemployment rate, the text output performs as well as, or better than, LME. On less common statistics, however, such as the proportion of mortgages that are at least 90 days delinquent (US counties) or the number of government employees (German districts), LME performs substantially better. The results are similar when choosing Pearson correlation as the performance metric, as shown in Figure AII in the Appendix.
Next we test whether we can improve the performance of LME by reducing the dimensionality of the embeddings using a PCA before training LME. For our two key datasets, US counties and US firms, Figures AIII and AIV in the appendix respectively show that LME usually performs best without applying a PCA.
We also compare the performance by layer and see that across the key datasets, embeddings from the 25th layer perform better on average than those from earlier or later layers (Figure AV in the appendix). However, the differences in performance are small with the exception of the lower performance when using layer 5, the earliest layer we examine. Table LABEL:tab:tab_cv_all presents the cross-validation results numerically for all of the datasets.
@@
Figure 2 plots the actual values against both the inferred values of LME (completion prompt, top row) and the text output (bottom row) on a few variables of the US counties dataset. The text output exactly recalls the population in a large number of counties, particularly for those counties with a large population. For the other variables, however, the text output is clustered on a few distinct values. [^1] Clustering of values in the text outputs has also been observed by Li et al. (2024) in their analysis of LLMs’ knowledge of regional statistics., [^2] Note that the values are not just clustered by US state. We typically observe several unique values within states as shown by Figure AVII which depicts the number of unique ground-truth and text output values for each state. By contrast, the distribution of the LME predictions, does not cluster but also does not reproduce the exact values.
For the US states dataset, we also evaluate the performance of the text output and LME within states. To obtain stable estimates we exclude states that have less than 50 counties. Figure AVI in the appendix compares the Spearman correlation between the ground truth and the text output and LME, respectively. Except for population, we observe a performance advantage of LME over the text output in most states for the remaining seven variables. [^3] We cannot show the variable proportion of mortgages being delinquent because it has a low coverage in the raw data, which is why we do not observe at least 50 non-missing observations in any state.

### v2

@@ Abstract
We investigate whether hidden states of large language models (LLMs) can be used to estimate and impute economic and financial statistics. Focusing on county-level (e.g. unemployment) and firm-level (e.g. total assets) variables, we show that a linear regression trained on the hidden states of open-source LLMs outperforms the models’ own text outputs. This indicates that internal representations encode richer economic information than is revealed directly in generated responses. A learning curve analysis shows that, in many cases, only a few dozen labelled examples suffice for training. We further propose a transfer learning method that improves estimation accuracy without requiring any labelled data for the target variable. Finally, we demonstrate the practical utility of hidden states in data imputation and super-resolution tasks.
@@ 1 Introduction
During training, generative large language models (LLMs) are exposed to vast amounts of information, including data relevant to economic modelling, such as geospatial statistics and firm-level financial metrics. If LLMs can effectively retrieve and utilise this knowledge, they could reduce dependence on external data sources that are time-consuming to access, clean, and merge, or that incur financial costs. Moreover, if LLMs accurately represent data, they could support downstream tasks like data imputation and outlier detection. In this study, we evaluate whether and how LLMs can be used for typical economic data processes.
@@
We test the usefulness of LLMs for data tasks with a large array of empirical analyses on regional economic data of the US, UK, EU, and Germany and financial data on US listed firms. First, we show that a regularised linear model learned on the hidden states often provides substantially better estimates of the economic and financial variables than the text output of the model, particularly for less common statistics. This result holds across open-source LLMs of varying sizes (1–70 billion parameters). A learning curve analysis finds that small samples of labelled data usually suffice to learn an accurate linear model on the embeddings. We also suggest a simple transfer learning algorithm that estimates a statistical variable without any labelled data for that variable. We achieve this by both learning from other labelled variables and using the text output of the LLMs as noisy labels on the variable of interest.
@@
Finally, we investigate the practical relevance of our findings for two data processes. Firstly, we show that we can exploit the embeddings to consistently improve the imputation of missing values in economic and financial variables, a common challenge in both research and industrial data pipelines. Secondly, we demonstrate that the hidden states of LLMs can be used for super-resolution tasks, inferring lower-level geographical statistics from high-level data, which can enable the estimation of statistics below the official reporting level. In both applications, we do not assume the LLM has seen the target statistics during training and can recall them by prompting; instead, we leverage the generalised, entity-level knowledge encoded in the model’s hidden states.
Our findings support the value of LLMs in data processing tasks, which warrants attention, given that tasks such as identifying data sources, extracting numeric information, merging data series, and handling missing values are time-consuming and error-prone. This is evident from the numerous data providers that sell packaged, cleaned, and well-structured datasets, even when a large proportion of the underlying data is publicly accessible online.
This study is structured as follows. Section 1.1 discusses the relevant literature, Section 2.1 describes the datasets and Section 2 outlines our empirical approach. The main results on the accuracy of our approach to estimate economic variables from embeddings are shown in Section 3. Section 4 presents the methodology and results for two data processing tasks using embeddings: imputation and super-resolution. Section 5 concludes.
@@ 1.1 Literature review
Our study is most closely related to papers that investigate how well LLMs can recall geospatial statistics. manvi2023geollm show that both GPT3.5 and fine-tuned Llama 2 can be prompted to reproduce geospatial information, such as population density, mean income, education, or house prices. The LLMs recall these variables more accurately than supervised machine learning methods, which inferred the information from other data sources such as satellite images of luminosity at night. Note that the authors only consider the text output of the LLM and do not exploit the hidden states as we do.
@@
Similarly, li2024can test how well LLMs can reproduce a large set of variables about cities and regions such as population density, life expectancy, average travel time, and number of patents. As in our study, this paper tests the accuracy of both the text output and a linear model on the hidden layer activations to estimate the statistics. However, they focus on a third approach: they ask the LLM to identify relevant features that can help to predict a variable of interest and then ask the LLM to generate values for these variables. After feeding these variables into a supervised ML model to predict the variable of interest, this approach performs on par with a linear model on the hidden states in 5-fold cross-validation. In contrast to our work, this paper focuses purely on how well the LLMs can recall statistics and does not consider transfer learning or downstream tasks.
@@
Our main empirical approach of learning a linear model on hidden states of a language model is known as linear probing in the literature (belinkov2022probing; alain2016understanding) and has been used to understand the inner workings and capabilities of LLMs. Our study builds in particular on the findings by gurnee2023language, showing that hidden states of LLMs linearly represent space and time. Concretely, a linear model trained on the hidden states of Llama 2 accurately infers the death of historical figures, the release year of artworks, or the geographical location of landmarks. godey2024scaling investigate the scaling laws of these results, showing that larger models represent space and time more accurately. Further, chen2023more perturb activations in the LLM to show that there is a causal connection between the representations and the text output.
@@
zhu2024language use linear probing to show that LLM embeddings represent the results of simple addition problems. While the exact number cannot be reconstructed using linear probes, the correlation between the actual number output and the inferred number (via the probes) is high. Linear probing is also used to gauge the truthfulness of statements (marks2023geometry; orgad2024llms; burger2024truth) and liu2024 show that hidden states can also be leveraged to estimate uncertainty of LLM responses in different tasks.
@@
Our work also relates to studies that use hidden states of LLMs for downstream tasks such as classification. Particularly, buckmann2024logistic showed how a linear model trained on hidden states of small LLMs after task-specific prompting can outperform GPT-4 when trained on only a few dozen samples per class (see also cho2023palp). Hidden states of LLMs have also proven to perform well as general text embeddings (lee2024nv; SFRAIResearch2024) as evidenced by their competitive performance on the Massive Text Embedding Benchmark (MTEB) (muennighoff2022mteb), which includes tasks such as clustering, re-ranking, retrieval and classification.
Our transfer learning strategy draws on the literature on deep learning from noisy labels (see song2022learning, for a review). Consistent with prior findings that deep networks can generalise beyond noisy supervision and achieve test accuracy above the quality of the training labels (rolnick2017deep; oyen2022robustness), we observe that our model outperforms the noisy labels derived from LLM outputs. Moreover, following evidence that neural networks learn clean signal early and only later memorise noise, we employ early stopping to curb overfitting (liu2020early).
@@
The literature on the use of LLMs for data imputation and super-resolution tasks – the two practical applications of our methodology that we explore in this study – is scarce. hayat2024claimdata fine-tuned Llama 2 to impute missing values given a textual description of the instance’s features. ding2024semantic use a similar approach for data imputation in recommender systems. By contrast, we do not rely on text output but use the embeddings of the LLM as additional features when employing standard imputation methods.
@@
While there exists a rich statistical and machine learning literature on the estimation of geospatial statistics at a granular level (tzavidis2018start; chi2022microestimates; aiken2022machine; viljanen2022machine), we are not aware of any study that explores the use of LLMs for this purpose.
@@ 2.1 Datasets
We use four public regional datasets covering US counties, EU regions, UK districts and German districts. Additionally, we use a dataset containing public financial information about US listed firms. The datasets are listed in Table 1. We assembled the datasets using different public sources, which are described in detail in Appendix A. All datasets are cross-sectional and are based on statistics from 2019, with the exception of the US firms dataset for which we obtain data from 2022.111The free tier of the Yahoo Finance API limits access to earlier history. Intentionally, all datasets predate the 2023 knowledge cut-offs of the LLMs we tested so that we can assess whether the LLMs can exactly recall statistics.
@@
Table 1: Datasets. Appendix A provides full documentation of datasets and sources.
Key variables that we use in the regional datasets are population, GDP per capita, unemployment rate, a measure of the income per capita, and a measure of life expectancy. Key variables we observe in the US firms dataset are total assets, market capitalisation, profitability metrics such as return on equity and return on assets. The full list of variables in all datasets are shown in Table LABEL:tab:tab_cv_all. For the super-resolution analysis in Section 4.2 we use additional data that we describe in that Section. We provide more extensive results for our main datasets, US counties and US listed firms, which contain a larger number of observations than the other datasets.
@@ 2.2 Experimental set-up
We exclusively use open-source LLMs, as, to the best of our knowledge, none of the providers of the leading proprietary models provide access to the hidden states of their generative models. Furthermore, we focus on smaller models because these are easier and cheaper to use. The models we test are listed in Table 2. In the main experiments, we consider three models of the Llama 3 family (1B, 8B and 70B parameters) and the Phi-3-mini (3.8B parameters) model. Apart from Llama 3 70B, we accessed the models through the transformers library (wolf2020transformers) on an NVIDIA Tesla V100 GPU. Due to its large size, we use a quantised version of Llama 3 70B on 2 NVIDIA T4 GPUs and obtain the embeddings with the llama-cpp-python library. To obtain this model’s text outputs, we use the non-quantised model hosted on replicate.com. More details on the computational requirements can be found in Appendix B.1.
We also consider a reasoning model: Qwen QwQ (32B parameters). To assess whether the reasoning paradigm improves the estimation of regional and financial statistics, we compare the model directly against Qwen 2.5 32B, which is the general-purpose LLM on which Qwen QwQ is based. To obtain the text output of both models, we use the instances hosted on groq.com. We obtain the embeddings of Qwen 2.5 32B from a quantised version using llama-cpp-python.
@@
When training a linear model on the embeddings, we also test a generic prompt, which does not vary with the variable of interest but just embeds the name of the entity and the year (e.g. “Orange County, California in 2019”).
@@ 2.3 Linear model on embeddings (LME)
We fit a ridge regression model on the hidden states of the prompt’s last token. Our baseline model has 32 layers, with a 4096-dimensional embedding vector. If not stated otherwise, we use the embeddings of the 25th layer because LME performs better on this layer than on others (Figure AV in the appendix).
To estimate the performance of LME, we employ grouped cross-validation on the regional datasets. Specifically, we ensure that the training and test sets do not share any values of the grouping variable. The grouping variable for each dataset is shown in Table 1. For US counties, for example, the grouping variable is US states. Our baseline results are based on 25 repeats of repeated 5-fold cross-validation but we also report learning curves, where, adhering to the grouping, we sample training sets of increasing size. [^2] We repeat the learning curve sampling procedure between 30 and 300 times, depending on the dataset size and training sample size.
We either train the linear model on the embeddings directly or on the first $k$ PCA components, with $k\in\{5,10,25,50,100,200\}$. We fit the PCA to all observations and not only on the training set. The regularisation parameter $\alpha$ is tuned over 50 logarithmically spaced candidates between $10^{-5}$ and $10^{5}$. The selected $\alpha$ minimises the mean squared error under 5-fold grouped cross-validation within the training set.
We denote the ground-truth values of the variable of interest by $y$, the embedding matrix by $E$, and the LLM’s textual output by $\overset{txt}{y}$.
@@ 2.4 Data transformations and performance metrics
To avoid a strong influence of outliers on the performance metrics and on the estimation of LME, we transform some of the dependent variables. Specifically, across all datasets, we applied a log transformation to non-negative variables with a skewed distribution. For skewed variables with negative values, we use a cubic transformation ($f(x)=sgn(x)\times|x|^{\frac{1}{3}}$). The transformations applied to each variable are reported in Table LABEL:tab:tab_cv_all.
When assessing the performance of the text output and LME, we test both models on exactly the same observations. The text outputs of the LLMs sometimes fall outside the range of expected values (e.g. unemployment rate >50>50%). In this case, we decide to remove the observations. Parsing numeric values from LLMs’ textual answers is challenging, due to the inconsistent format and inconsistent use of units. We made considerable effort to correctly parse the text output and conducted comprehensive manual checks on all variables to ensure the accuracy of our parsing approach. Our parsing approach is described in detail in Appendix B.3.
@@ 3 Using LLM embeddings to infer statistics
Figure 1: Cross-validation performance. The variable labels are shortened. The full variable names used when querying the LLM are shown in Table LABEL:tab:tab_cv_all.
@@ 3.1 Cross-validation
Figure 1 compares the performance (Spearman correlation) of the text output to the baseline LME learned on the embeddings of the completion prompt and the generic prompt. The variables are ordered by decreasing performance of the text output. [^3] See Table LABEL:tab:tab_cv_all for a tabular representation of the results including a measure of uncertainty, which we omitted from these charts to improve legibility. LME based on the completion prompt performs better than LME on the generic prompt in all datasets except the US firms data, where both perform equally well on most variables. We therefore focus on the completion prompt in the following analyses.
LME beats the text output most of the time. The performance advantage of LME increases with decreasing performance of the text output. On more common variables, such as population, or unemployment rate, the text output performs as well as, or better than, LME. On less common statistics, however, such as the proportion of mortgages that are at least 90 days delinquent (US counties) or the number of government employees (German districts), LME performs substantially better. The results are similar when choosing Pearson correlation as the performance metric, as shown in Table LABEL:tab:tab_cv_all and Figure AII in the Appendix.
@@
Figure 2 plots the actual values against both the inferred values of LME (completion prompt, top row) and the text output (bottom row) on a few variables of the US counties dataset. The text output exactly recalls the population in a large number of counties, particularly for those counties with a large population. For the other variables, however, the text output is clustered on a few distinct values. [^4] Clustering of values in the text outputs has also been observed by li2024can in their analysis of LLMs’ knowledge of regional statistics., [^5] Note that the values are not just clustered by US state. We typically observe several unique values within states as shown by Figure AVII which depicts the number of unique ground-truth and text output values for each state. By contrast, the distribution of LME predictions does not cluster but also does not reproduce the exact values.
For the US states dataset, we also evaluate the performance of the text output and LME within states. To obtain stable estimates we exclude states that have less than 50 counties. Figure AVI in the appendix compares the Spearman correlation between the ground truth and the text output and LME, respectively. Except for the population variable, we observe a performance advantage of LME over the text output in most states. [^6] We cannot show the variable proportion of mortgages being delinquent because it has a low coverage in the raw data, which is why we do not observe at least 50 non-missing observations in any state.
3.2 Robustness
We test whether we can improve the performance of LME by reducing the dimensionality of the embeddings using a PCA before training LME. For our two key datasets, US counties and US firms, Figures AIII and AIV in the appendix respectively show that LME usually performs best without applying PCA.
We also compare the performance by layer and see that across the key datasets, embeddings from the 25th layer perform better on average than those from earlier or later layers (Figure AV in the appendix). However, the differences in performance are small with the exception of the lower performance when using layer 5, the earliest layer we examine.